### v1

# Semantic Latent Motion for Portrait Video Generation

Abstract
Recent advancements in portrait video generation have been noteworthy. However, existing methods rely heavily on human priors and pre-trained generation models, which may introduce unrealistic motion and lead to inefficient inference. To address these challenges, we propose Semantic Latent Motion (SeMo), a compact and expressive motion representation. Leveraging this representation, our approach achieve both high-quality visual results and efficient inference. SeMo follows an effective three-step framework: Abstraction, Reasoning, and Generation. First, in the Abstraction step, we use a carefully designed Mask Motion Encoder to compress the subject’s motion state into a compact and abstract latent motion (1D token). Second, in the Reasoning step, long-term modeling and efficient reasoning are performed in this latent space to generate motion sequences. Finally, in the Generation step, the motion dynamics serve as conditional information to guide the generation model in synthesizing realistic transitions from reference frames to target frames. Thanks to the compact and descriptive nature of Semantic Latent Motion, our method enables real-time video generation with highly realistic motion. User studies demonstrate that our approach surpasses state-of-the-art models with an 81% win rate in realism. Extensive experiments further highlight its strong compression capability, reconstruction quality, and generative potential. Moreover, its fully self-supervised nature suggests promising applications in broader video generation tasks.
1 Introduction
Portrait video generation has gained significant attention due to its numerous applications in gaming, film, and education. With the development of AI generation technologies, there is an increasing demand for more realistic animation effects and faster generation speeds.
Figure 1: Semantic Latent Motion is a compact and expressive representation of the character’s current state. The motion dynamics in the semantic latent space serve as a powerful guiding condition, steering the generative model from the reference image to the target image. Leveraging the compactness of latent motion, we can model long sequences and perform more efficient resoning.
Figure 2: Overview of SeMo. The proposed framework follows a three-step process (abstraction, reasoning, generation) implemented in two training stages. In the first stage, we train a Motion AutoEncoder for abstraction and generation. During the abstraction step, image features are subjected to random masking and compression to obtain a compact and abstract latent motion. In the generation step, latent motion is used as a condition to guide the diffusion decoder in generating the target image from the reference image. In the second stage, we train a Motion Generator for reasoning, which generates motion sequences from audio.
Existing approaches mainly follow two paths. The first path integrates human prior knowledge of motion representation, such as sequences of 3DMM, pose, landmarks, etc. as expressions of motion [9, 4], combined with a generative renderer to produce the final video frames. The second path utilizes pre-trained video generation models, using their strong video generation capabilities and conditional control of cross-attention to learn the mapping from audio to video frames [6, 4, 41]. Many methods combine both of these paths. However, in the first path, human prior knowledge relies on domain-specific heuristics, often failing to capture detailed motions, which results in human videos that lack naturalness and coherence. Moreover, this framework is not adaptable to multiple scenarios (for example, it cannot be used for general video generation, as it is impractical to extract pose sequences for all objects in the scene). In the second path, the complex spatiotemporal attention modeling and multi-step denoising in diffusion models prevent real-time generation and long-context modeling.
We observe that these motion representations ( 3DMM, pose, landmarks ) all express a high-level semantic:“ the positions and attributes of meaningful parts.” They are all abstractions of human knowledge about facial dynamics and are limited by design biases. For instance, we place landmarks on faces because we believe that changes in the positions of the noise, eyes, and mouth corners are crucial for facial dynamics. Similarly, we use skeletons because we believe that skeletal structures are important for body movements. However, these representations are limited by human abstracting capabilities and cannot capture detailed motion. We ask: Can deep learning models abstract a semantic latent representation that better captures detailed face motions by self-supervised learning? We will introduce our motivation from the following two aspects.
1) Representing video motion in a compact and expressive space. Modeling video motion presents significant challenges. Fortunately, facial videos exhibit more constrained motion ranges and patterns compared to general videos. Previous works have demonstrated that models can synthesize realistic images by extracting detailed information from reference images and semantic information from prior knowledge. Human-crafted priors (e.g., facial landmarks) represent compact and semantic motion representations abstracted through domain knowledge. If humans can achieve this, deep learning models could potentially learn superior representations through self-supervision – capturing more compact and detail-preserving motion characteristics. This motivates our exploration of learning semantic latent motion representations that surpass human-designed priors in expressiveness.
Figure 3: Given an audio clip and a reference image, SeMo can generate lifelike video. From top to bottom, we highlight several key features of our model: a. The generated characters exhibit natural head movements and blinking while speaking. b. They display rich motion dynamics with substantial amplitude. c. Model can adapt to different poses. d. Model generalize well across various datasets. e. The generated videos feature remarkably natural lighting and shadow variations.
2) Employing self-supervised semantic representations to guide generation. Current generative models [5, 17, 1, 2, 43, 23, 46, 13] predominantly rely on human-abstracted conditions like text prompts or landmarks. While effective, these representations are inherently limited by human abstraction capabilities and domain-specific biases, often failing to capture subtle motion details. Recent works like RCG [18] and StoryDiffusion[52] have demonstrated that self-supervised latent features can serve as superior generation conditions compared to manual priors. Building on this insight, we propose using learned semantic latent motion as conditional signals for video generation, enabling more precise and detailed motion control.
Motivated by these insights, we design a universal motion representation framework coupled with a simple yet effective video generation pipeline. Our approach consists of two key components: First, we train a Motion Autoencoder, where an encoder compresses each facial video frame into a compact latent motion representation (1D token) using a random masking strategy. The diffusion-based decoder then reconstructs the target frames using reference images as inputs and semantic latent motion as conditional signals. Once the Motion Autoencoder is trained, we extract latent motion representations for downstream tasks. For portrait video generation, we train a diffusion-based generator that synthesizes latent motion sequences directly from audio inputs. The compact latent space enables efficient long-term temporal modeling and preserves comprehensive event information. During inference, we first generate motion sequences from audio and reference images, then synthesize video frames using the Motion Autoencoder’s decoder.
The rigorous reconstruction task encourages our latent motion representation to develop strong descriptive power, effectively capturing essential information within an extremely compact space. Leveraging this representation, our method achieves highly natural and vivid animations, outperforming state-of-the-art methods with an FID score of 26.8 (compared to the 27.4 baseline) and achieving an 81% win rate in realism in a user study.
Beyond these results, we further observe that even when 90% of the video frame content (input to the motion encoder) is masked, our Motion AutoEncoder can still generate coherent and smooth videos. This reveals two critical insights: 1) High-level semantic information sufficiently guides video generation, 2) Video motion modeling requires less information proportion than image generation (contrasting 75% masking ratio in MAE [14]). Moreover, as a general framework, our Motion AutoEncoder demonstrates strong adaptability to generic videos containing limited subjects and simple motions, such as object rotations.
2 Related Works
Motion-Based Video Codec.
Motion-driven video synthesis techniques typically generate target frames by combining reference frames with motion representations. FOMM [30] established a baseline through sparse keypoint detection and affine transformation estimation to derive dense motion fields. LIA [39] advanced this paradigm by decomposing latent vectors and employing optical flow-based warping operations. MRAA [31] extended the framework by separately predicting foreground region motions and background displacements, subsequently rendering through pixel-wise flow manipulation. IMF [10] introduced a fully self-supervised approach that encodes reference and target frames into compact latent tokens. This method progressively reconstructs target frame pixels through a coarse-to-fine process that integrates reference image features at multiple granularity levels. These developments collectively demonstrate the evolution from explicit motion parameterization towards more implicit, learned representations that better preserve spatial-temporal coherence.
Portrait Video Generation.
The portrait generation task has garnered significant attention in recent years due to its broad applications and has witnessed rapid advancements. Wave2Lip [25] pioneered the use of GAN networks combined with lower-face masking to directly synthesize video frames from reference images and audio inputs. Subsequent works including VividTalk [33], SadTalker [48], and Moda [22] have leveraged 3D morphable priors coupled with neural rendering techniques to generate target video frames. Notably, VASA-1 [42] and Gaia [15] introduced hybrid architectures that encode both 3D prior knowledge and appearance features into a latent space, establishing a more universal and compact motion representation framework. Next, diffusion-based generators are employed to produce motion sequences conditioned on audio inputs [34, 19, 32, 49]. Recent innovations like EMO [34] and Hallo2 [6] have further enhanced visual quality through the integration of pre-trained video generation foundation models with specialized reference networks. AniTalker [21] presents a two-stage self-supervised learning paradigm. The overarching trend in this domain demonstrates a clear shift from hand-crafted priors towards data-driven approaches that enable models to capture finer details and enhance the naturalness of synthesized visuals.
Self-supervised Representation as Guidance.
A large amount of self-supervised learning work  [14, 28, 35, 11, 26, 44, 37] has utilized encoded features for various downstream visual understanding tasks. Self-supervised representations have been proven to possess rich semantics and can serve as powerful guidance signals. In RCG  [18], the quality of image generation was significantly improved by introducing features obtained through MoCo’s [3] self-supervised learning framework. Meanwhile, Story Diffusion [52] obtains a continuous guidance signal by performing linear interpolation in the latent space of CLIP features, thereby controlling the video generation model to produce more coherent content.
3 Method
Figure 4: Illustration of Motion AutoEncoder
Figure 5: An example of video reconstruction: In the example above, we demonstrate that SeMo can not only accurately model facial movements but also capture environmental changes. With a token size of 1, semantic latent motion is not only more compact than the motion in other works but also captures more comprehensive information.
Given a reference image and its corresponding audio, our goal is to generate a dynamic video with synchronized motion and sound. To achieve this, we propose a two-stage self-supervised learning framework. In the first stage, the objective is to extract a compact, high-level semantic motion $\mathbf{M}\in\mathbb{R}^{l\times d}$ for each frame of the image. We first use learnable tokens and random masking to compress the motion into an abstract representation, and then utilize the reference image and the motion to reconstruct the target image. Once the Motion AutoEncoder is trained, it is frozen. In the second stage, we train a Motion Generator. For each frame, we first use the motion encoder to extract the motion and then generate the target motion based on the audio and the corresponding motion of the reference image. The details of these two stages will be described below.
3.1 Motion AutoEncoder
Motion Encoder. Our motion encoder design is motivated by the observation that facial movements - akin to 3D Morphable Models (3DMM) and keypoint trajectories - can be effectively represented in a highly compact space. We propose a novel Masked Motion Encoder architecture that integrates dual strategies of random masking and latent space compression to derive semantically rich latent motion.
Specifically, we first compress input frames $\mathbf{I}\in\mathbb{R}^{H\times W\times C}$ into latent representations $\mathbf{X}\in\mathbb{R}^{h\times w\times c}$ using a pre-trained stable diffusion VAE encoder [29]. Following ViT [8], we embed each latent feature into non-overlapping patches $\mathbf{Z}\in\mathbb{R}^{t\times d}$, where $t=\frac{H}{p}\times\frac{W}{p}$ denotes the token count of each image, $p$ denotes the patch size, and apply random masking with ratios uniformly sampled from 0% to 90%. The masking ratios are carefully designed, low mask ratios are used to encourage the model to retain more complete information, while high mask ratios aim to enhance the model’s ability to abstract. Once the masking is complete, we discard the masked tokens and retain the unmasked tokens, arranging them into a sequence in their original order. As shown in Fig. 10, we demonstrate that even with 90% of the content masked, the model is still able to generate acceptable results.
$\small\mathbf{M}=\operatorname{Enc}(\mathbf{Z^{m}}\oplus\mathbf{T})$ (1)
where $\oplus$ denotes concatenation, $\operatorname{Enc}$ denotes Motion Encoder, $\mathbf{Z^{m}}$ denotes masked image patches.
TikTok [45] has demonstrated that image features can be represented as 1D sequence representations. Therefore, we designed a 1D learnable token sequence, which is adjustable in length and can become very compact. As shown in Fig. 4, we concatenate the learnable tokens with the image features and process them through ViT layers. Ultimately, latent motion $\mathbf{M}$ is output at the position corresponding to the learnable tokens. Here, the number of motion tokens and the token dimension size are consistent with the learnable tokens. Note that, unless otherwise specified, our latent motion is set to 1 token with a dimension of 512. Tab. 5 Tab. 6 varies number and dimension of tokens.
Diffusion Decoder. At the stage, our objective is to use the reference image as a reference for detailed information, with motion serving as the semantic condition, guiding a generative model to fill in missing details. Our decoder is a standard DiT [24] model. Specifically, we randomly select two sets of images from the data, obtaining reference image $\mathbf{Z_{r}}$ and target image feature $\mathbf{Z_{t}}$. We apply the same masking ratio to image features and obtain reference motion $\mathbf{M_{r}}$ and target motion $\mathbf{M_{t}}$ through the motion encoder. Following RectifiedFlow[20], We add noise to the unmasked target image feature Eq. 2, and concatenate noised target image feature $\mathbf{Z_{t}^{n}}$ with the $\mathbf{Z_{r}}$, $\mathbf{M_{r}}$ and $\mathbf{M_{t}}$ Eq. 3. Finally, the DiT decoder outputs the predicted noise velocity $\hat{\mathbf{V}}$. We refer to the module described above as the Transfer Block.
$\small\mathbf{Z_{t}^{n}}=(1-t)\mathbf{Z_{t}}+t\epsilon$ (2)
$\small\hat{\mathbf{V}}=\operatorname{Dec}(\mathbf{Z_{r}}\oplus\mathbf{Z_{t}^{n}}\oplus\mathbf{M_{r}}\oplus\mathbf{M_{t}})$ (3)
Where $\epsilon$ denotes Gaussian Noise sampled from standard normal distribution, $t\in[0,1]$ denotes the intensity of the noise, $\mathbf{Z_{t}^{n}}$ denotes noisy target image feature.
The above module enables the transfer from a reference image to a target image. We can extend this framework to transfer from the reference image to target video. Specifically, we randomly select one reference frame and a sequence of $f$ consecutive video frames. After applying Transfer Block to each frame, we introduce a Temporal Align Block. This is because we observed that treating each frame independently for denoising results in a lack of smooth transitions in the details. In Temporal Align Block, We flatten noised target image feature $\mathbf{Z^{n}_{t}}\in\mathbb{R}^{f\times t\times d}$ to $\mathbf{Z^{n}_{t}}\in\mathbb{R}^{b\times d}$, where $f$ denotes number of frames, $b=f\times t$. After transformation, we apply self-attention for temporal modeling, The complexity of this block is $o(f)$, which does not lead to significant computational overhead.
3.2 Motion Generator
In this module, our objective is to train a motion generator that can generate a corresponding motion sequence based on a previous motion sequence, a randomly selected reference motion, and an audio sequence. Our generator is a standard Diffusion Transformer (DiT). We first apply a 0–20% mask Eq. 4 to the previous image feature before extracting the reference motions $\mathbf{M_{p}^{m}}$ and extract target motions $\mathbf{M_{t}}$ without mask Eq. 5, we add noise to the target motions Eq. 6, these are concatenated along the time dimension and processed through self-attention layers. Following this, we use two cross-attention layers: one with the audio feature $\mathbf{A}$ and the other with the reference motion $\mathbf{M_{r}}$, note that the audio feature is embeded by Whisper [27]. In Fig. 9, we demonstrate that applying a mask to the previous image enhances the model’s generalization ability.
$\small\mathbf{M_{p}^{m}}=\operatorname{Enc}(\mathbf{Z^{m}_{p}}\oplus\mathbf{T})$ (4)
$\small\mathbf{M_{t}}=\operatorname{Enc}(\mathbf{Z_{t}}\oplus\mathbf{T})$ (5)
$\small\mathbf{M_{t}^{n}}=(1-t)\mathbf{M_{t}}+t\epsilon$ (6)
$\small\hat{\mathbf{V}}=\operatorname{G}\left(\mathbf{M_{p}^{m}},\mathbf{M_{t}^{n}},\mathbf{M_{r}},\mathbf{A}\right)$ (7)
3.3 Training Objective
In the first stage Eq. 8, the goal of the motion autoencoder is to reconstruct the target images based on the reference image and corresponding motion. In the second stage Eq. 9, the model is optimized to generate motion sequence based on audio condition. In both stage, the model predicts the noise velocity, and we optimize the mean squared distance between the predicted velocity $v_{\theta}$ and the true velocity $v$.
$\mathcal{L}_{1}=\mathbb{E}_{t,z_{r},z_{t},\epsilon}\left[\left\|(z_{t}-\epsilon)-v_{\theta}\left(z_{r},z_{t},t\right)\right\|^{2}\right]$ (8)
$\mathcal{L}_{2}=\mathbb{E}_{t,m_{r},m_{t},m_{p},\epsilon}\left[\left\|(m_{t}-\epsilon)-v_{\theta}\left(m_{r},m_{t},m_{p},t\right)\right\|^{2}\right]$ (9)
4 Experiment
4.1 Experimental Settings
Figure 6: Qualitative comparison with existing approaches. The results demonstrate that our model effectively generates complex, realistic, and natural motions across various audio conditions.
Table 1: Model Capacity of various methods.
Method Unsupervised Long-term Modeling Real-time Generation Generic Generation Pipeline
Aniportrait [41]
$\times$
$\times$
$\times$
$\times$
Hallo2 [6]
$\times$
$\times$
$\times$
$\times$
SadTalker [48]
$\times$
$\times$
Anitalker [21]
$\times$
Echomimic [4]
$\times$
$\times$
$\times$
$\times$
SeMo (Ours)
Table 2: Motion-based video reconstruction comparison of various methods on the HDTF, DH-FaceVid-1K, and MEAD dataset.
Method HDTF DH-FaceVid-1K MEAD
[…]
FOMM[30] 24.690 188.797 0.064 25.572 0.816 73.484 30.052 236.578 0.061 25.542 0.844 60.701 41.541 166.950 0.040 31.131 0.917 60.267
MRAA[31] 23.004 155.53 0.046 29.411 0.891 73.653 22.258 216.034 0.044 30.295 0.913 60.481 35.378 167.186 0.030 34.245 0.950 50.558
TPSMM[51] 26.556 138.88 0.041 29.51 0.901 58.478 24.697 215.362 0.042 29.625 0.917 54.784 34.358 136.651 0.028 34.805 0.956 69.224
LIA[39] 26.366 186.307 0.067 26.624 0.840 66.068 33.409 238.101 0.067 26.876 0.861 54.488 47.152 170.724 0.048 31.124 0.921 45.276
Liveportrait[12] 22.469 137.935 0.068 24.810 0.802 59.955 26.502 170.360 0.059 25.681 0.839 52.026 56.700 88.304 0.038 31.661 0.923 44.817
SeMo (Ours) 21.431 136.738 0.043 29.12 0.912 57.849 22.045 112.166 0.036 32.132 0.914 51.437 23.546 70.975 0.019 36.233 0.949 44.517
Datasets. For training the Motion AutoEncoder, we utilized three datasets: DH-FaceVid-1K [7], HDTF [50], and MEAD [36]. Videos were resized and center-cropped to 256×256 resolution, excluding those failing face detection. Training samples were equally drawn from all three datasets. For training the Motion Generator, we employed two datasets: HDTF and DH-FaceVid-1K. We removed samples with excessively noisy or unclear audio and encoded the audio using the pre-trained model Whisper [27].
Implementation Details. All two-stage experiments were trained on 8 NVIDIA A800 GPUs. In Stage 1, we first trained an image-to-image version of the Motion AutoEncoder. Subsequently, we utilized the pre-trained weights to further train it as an image-to-video version. In the main experiments, we maintained the motion representation as 1 token with dimension 512. In Stage 2, we trained the motion generator, generating 16 consecutive motion tokens in a single forward pass. A consistent learning rate of 1e-4 was applied across all experiments.
Evaluation Metrics. We evaluate our experiments very rigorously using various metrics. In the video reconstruction phase, we use FID [16], FVD [38], LPIPS [47], SSIM [40] and PSNR. In the audio-driven portrait video generation phase, we use FID, FVD, Sync-C and Sync-D.
Baseline. In the motion-based video reconstruction phase, we compared MRAA [31], FOMM [30], Lia [39], Liveportrait [12], TPSMM [51]. In the portrait video generation phase, we compared Aniportrait [41], Anitalker [21], Echomimic [4], Hallo2 [6] , SadTalker [48]
4.2 Comparisons.
Motion-based Video Reconstruction. Motion-based video reconstruction refers to an algorithm that generates target frames based on the reference image and motion. In the experiment, Tab. 2 we directly applied their pretrained weights in the face reconstruction task. We observed that our method outperformed theirs in terms of metrics. Additionally, our approach not only modeled the motion of the face but also modeled the motion in the surrounding areas of the face, making the entire video look more realistic.
Audio-driven Video Generation. Our generation framework achieves remarkably natural motion effects. As shown in Fig. 5 , when processing utterances like ”Yes!”, existing methods only produce limited lip movements, while our approach naturally integrates lip pursing with spontaneous eye blinks to deliver realistic contemplative expressions. Fig. 3 highlights our framework’s robustness across diverse facial poses and camera distances, generating substantial yet plausible motion variations. Notably, our method effectively captures authentic lighting interactions (e.g., dynamic shadow transitions during head movements), significantly enhancing visual realism. Quantitative evaluations in Tab. 4 Tab. 3 confirm our superiority over SOTA methods, with particularly significant improvements in FVD.
Tab. 1 demonstrates the strong capabilities of the model. The framework’s scalability stems from our compact latent motion representation, enabling long-term temporal modeling for holistic event understanding. Through RectifiedFlow’s reflow distillation, we achieve real-time synthesis with minimal sampling steps. As a prior-free and unsupervised architecture, our framework demonstrates strong cross-domain adaptability, enabling efficient application to novel scenarios without architectural modifications.
Table 3: Audio-driven video generation comparison of various methods on the DH-FaceVid-1K dataset.
Method DH-FaceVid-1K
[…]
AniPortrait [41] 44.807 484.861 0.168 19.103 0.671
Hallo2 [6] 61.803 698.668 0.155 20.016 0.658
SadTalker [48] 85.447 610.627 0.328 14.382 0.538
Anitalker [21] 56.322 556.607 0.289 15.064 0.572
Echomimic [4] 69.296 691.534 0.327 14.218 0.527
SeMo (Ours) 27.667 206.523 0.095 22.954 0.768
Table 4: Audio-driven video generation comparison of various methods on the HDTF dataset.
Method HDTF
[…]
AniPortrait [41] 33.139 422.644 0.138 20.042 0.672 5.817 10.881
Hallo2 [6] 27.447 276.564 0.105 21.692 0.713 6.284 8.499
SadTalker [48] 93.744 548.635 0.225 14.493 0.437 5.439 9.940
Anitalker [21] 41.034 430.082 0.208 17.434 0.588 5.927 8.724
Echomimic [4] 70.768 591.313 0.311 14.920 0.452 6.030 10.462
SeMo (Ours) 26.860 218.414 0.111 21.985 0.739 6.147 8.574
User Study. We randomly selected 10 images from HDTF and DH-FaceVid-1K, and generated 20 videos using 2 random audio clips. We invited 20 participants to choose the best model from 4 different perspectives. We significantly outperform competitors in realism and match the state-of-the-art (SOTA) in lip sync, smoothness, and overall quality.
4.3 Ablation Study
Dimension and Token Number. In Tab. 5 Tab. 6, we compared the impact of latent motion token dimensions and token numbers on model performance. For the token dimension, we fixed the number at 1 and varied its size, finding that even with a dimension as small as 32, the model could still generate reasonable images. However, smaller dimensions resulted in reduced detailed information, requiring the diffusion model to compensate, which introduced some randomness in the output, though the layout remained consistent. For the token number, we fixed the motion representation space at 512 floats and found that increasing the number of tokens improved performance. This is due to the increased parameters used by the model to analyze motion, enabling more accurate mapping from semantic tokens to low-level pixels.
Table 5: Ablation studies of the latent Motion’s channel dimension on the HDTF dataset.
Token Dimension HDTF
[…]
T1D32 29.877 237.652 0.068 26.916 0.824
T1D64 26.174 214.236 0.066 27.520 0.835
T1D128 24.467 192.702 0.053 27.587 0.837
T1D256 22.629 157.235 0.048 28.430 0.865
T1D512 21.431 136.738 0.043 29.12 0.871
T1D768 21.424 135.678 0.042 29.385 0.872
T1D1024 21.369 133.547 0.039 29.430 0.877
Table 6: Ablation studies of the latent Motion’s token length on the HDTF dataset.
Token Number HDTF
[…]
T1D512 21.431 136.738 0.043 29.120 0.871
T2D256 21.082 133.529 0.041 29.561 0.873
T4D128 20.561 130.061 0.040 29.992 0.880
T8D64 20.138 128.390 0.038 30.180 0.883
Figure 7: Unmasked Motion AutoEncoder training retains details in latent motion, while masking reduces them.
Mask Strategy. In Fig. 10 We compared the impact of using a mask strategy during the training of both the Motion AutoEncoder and Motion Generator. In training the Motion AutoEncoder, we applied a mask to both the reference and target image features, sampling the mask ratio from 0% to 90%. While reconstruction metrics showed similar results with or without the mask, we found that the version without the mask included more detailed motion information, aiding the diffusion decoder. However, the mask did not significantly affect reconstruction outcomes, and even with the mask, good results were achieved. When replacing the reference image with a different person, the version without the mask captured irrelevant details like background and clothing, while the masked version focused more on motion-related details from the reference image, aligning with our goal of abstracting motion while maintaining reconstruction quality. In training the Motion Generator, we applied a mask of 0-20% to reduce the generator’s reliance on detailed information from the previous motion and encourage it to extract details from the reference motion. Without the mask, the model struggled to generalize to other datasets, but with the mask, its generation capabilities improved and it became more adaptable to new datasets.
Table 7: Ablation studies of the Mask Strategy
Method HDTF
[…]
w/ mask 21.431 136.738 0.043 29.120 0.871
w/o mask 21.326 136.229 0.041 29.061 0.876
Figure 8: User study. The numbers in the figure indicate the win rate for each dimension.
Figure 9: Unmasked motion encoding causes overfitting to details, compromising cross-dataset generalization.
5 Discussion and Limitation
Figure 10: Motion encoding in SeMo requires minimal information, achieving good results even when 90% of the data is masked.
Figure 11: SeMo can be adapted to general video scenes.
Motion encoding requires minimal information. As demonstrated in the MAE [14], model needs very minimal to generate meaningful visual content. However, in the context of motion encoding, even less information is required. The core difference between motion encoding and image encoding lies in the fact that motion encoding has a reference image. This makes much of the detailed information in the target image unnecessary. Previous methods did not make good use of the information in the reference image. In the example Fig. 10, we observe that even 90% of the image is masked, the original video can still be reconstructed.
Transfer to General Scenarios. We investigated whether Motion AutoEncoder is applicable to general videos. In Fig. 11, We found that in single-subject motion scenarios, Motion AutoEncoder performs well in reconstruction, even effectively learning hand movements. However, in complex multi-subject motion scenarios, noticeable distortions occur. We reproduced IMF and train them on WebVid.
Abstract Resoning and Detail Generation. Current video generation models, like Sora, perform temporal reasoning and pixel generation simultaneously, neglecting redundancy in motion representation. Consider how humans think – we don’t fully visualize the next state, but rather abstract conceptual representations for efficient reasoning. Video generation models could adopt similar principles: by constructing a more compact informational space through abstraction, we enable efficient inference. Subsequently, the abstract information guides the generative model in pixel synthesis, making prediction more efficient and manageable.
Limitation. We found that the generated results have blurry teeth, and high-frequency areas like beards tend to flicker. We believe this is caused by the factor that: we resize data to 256x256 and compression into the latent space via SD VAE, there was significant information loss. However, our method is scalable, and we can easily extend it to higher resolutions, by maintaining the current approach.
6 Conclusion
In this work, we have proposed a fully self-supervised framework for portrait video generation, structured around three key steps: Abstraction, Reasoning, and Generation. First, we employ a masking strategy and a video reconstruction task to compress motion into a compact, abstract space (1D token). Second, within this space, we perform long-term modeling and efficient reasoning. Finally, the inferred motion serves as a conditioning signal to guide the diffusion model in generating fine-grained details. Furthermore, our findings indicate that motion can be further compressed, enabling a more compact and efficient abstract representation. Beyond portrait video generation, our pipeline is broadly applicable to general video generation tasks. In future research, we aim to explore its feasibility across a wider range of scenarios.